# An Information Bottleneck Approach for Markov Model Construction

### v1

@@ IV Discussion
In terms of SPIB hyperparameters, our observations indicate robust performance across a broad range of hyperparameter values (refer to results in the SI). However, for achieving an optimal MSMs performance, we recommend employing the GMRQ score and cross-validation for hyperparameter tuning. The GMRQ measures the ability of the model to capture the slowest dynamics while the cross-validation procedure allows for a robust quantification of the model’s performance on new data. Our recommended tuning process begins with fixing a large lag time $\Delta t$ to derive a small number of macrostates while subsequently tuning the other hyperparameters in a sequential manner. Once the optimization of all other hyperparameters is complete, the next step involves scanning the lag time $\Delta t$ for SPIB. This step will generate varying numbers of converged states for different $\Delta t$. To finally determine the lag time $\Delta t$, one can either select the plateau (if there is a clear timescale separation) or specify $\Delta t$ based on the time resolution that we care about in a dynamical system. Alternatively, one may choose the lag time yielding the desired number of states for MSMs construction. For various conformational changes of biological molecular systems, previous studies elucidate their free energy landscape is hierarchical.[65, 66] A smaller $\Delta t$ may yield higher-resolution MSMs, identifying numerous free energy minima separated by small barriers as metastable states, whereas a larger $\Delta t$ may result in lower-resolution MSMs, detecting fewer metastable states encompassing multiple local free energy minima. SPIB provides great flexibility for users to select the most suitable resolution for MSMs to investigate various biological phenomena.

### v2

@@ IV Discussion
In terms of SPIB hyperparameters, our observations indicate robust performance across a broad range of hyperparameter values (refer to results in the SI). However, for achieving an optimal MSMs performance, we recommend employing the GMRQ score and cross-validation for hyperparameter tuning. The GMRQ measures the ability of the model to capture the slowest dynamics, while the cross-validation procedure allows for a robust quantification of the model’s performance on new data. Our recommended tuning process begins with fixing a large lag time $\Delta t$ to derive a small number of macrostates while subsequently tuning the other hyperparameters in a sequential manner. Once the optimization of all other hyperparameters is complete, the next step involves scanning the lag time $\Delta t$ for SPIB. This step will generate varying numbers of converged states for different $\Delta t$. To finally determine the lag time $\Delta t$, one can either select the plateau (if there is a clear timescale separation) or specify $\Delta t$ based on the time resolution that we care about in a dynamical system. Alternatively, one may choose the lag time yielding the desired number of states for MSM construction. For various conformational changes of biological molecular systems, previous studies elucidate their free energy landscape is hierarchical.[65, 66] A smaller $\Delta t$ may yield higher-resolution MSMs, identifying numerous free energy minima separated by small barriers as metastable states, whereas a larger $\Delta t$ may result in lower-resolution MSMs, detecting fewer metastable states encompassing multiple local free energy minima. SPIB provides great flexibility for users to select the most suitable resolution for MSMs to investigate various biological phenomena.
We note that, while the SPIB models shown here typically show good performance regarding the GMRQ score, the SPIB loss function (eq. LABEL:eq:SPIB_obj) is not equivalent to the GMRQ or VAMP-based scores, though Both VAMP-based scores and SPIB loss aim to approximate the propagator of underlying dynamics. The GMRQ score used in VAMP-based approaches relies on maximizing the sum of the top n eigenvalues of transition probability matrix constructed on the coarse-grained subspace of the relevant, slow dynamics. In contrast, the SPIB loss function is the evidence lower bound to the maximum likelihood estimate of the state transition density. It comprises a reconstruction term and a regularization term. In the limit that the hyper-parameter $\beta$ governing the regularization term in the SPIB loss function approaches zero, the SPIB loss function focuses solely on minimizing the error of propagator approximation. Consequently, optimizing the SPIB loss in this scenario may yield outcomes akin to those derived from the GMRQ score. However, it’s worth noting that a key distinction from VAMP-based methods remains: SPIB focuses on slow transitions with significant Boltzmann weights, whereas VAMP-based approaches concentrate solely on the slowest processes without considering their weights. This unique aspect allows SPIB to consistently learn highly populated metastable states, resulting in higher entropy scores in the systems tested in this study.
In this work, to maintain consistency with prevailing practices in the MSM literature, we chose to conduct hyperparameter selection based on the GMRQ rather than the SPIB loss function. Nevertheless, recognizing both the connections and distinctions between the SPIB loss and VAMP-based scores, we acknowledge the potential of exploring the direct use of the SPIB loss for cross-validation as a valuable avenue for future research.
The combination of the SPIB loss function, which incorporates a heuristic for state metastability, with the flexible use of lag time $\Delta$t allows for the construction of multi-resolution MSMs that perform well over the range of evaluation metrics provided in Tables 1 and 2. Furthermore, the ability of the SPIB to learn adaptively the number of metastable states on-the-fly and as a function of lagtime differentiate the SPIB models from the analogous VMAPnet models.
In addition, the SPIB consistently learns a low-dimensional latent space on-the-fly, effectively embedding the transitions between multiple metastable states of interest. This contrasts with existing dimension reduction methods, such as tICA and PCA, which typically fall short in capturing transitions between multiple metastable states within a 2D space (see SI).
Overall, we believe this one-shot nature of the SPIB, where a latent space for description of the slow dynamics and kinetic modeling is built simultaneously, provides a low dimensional embedding of the metastable states of interest and is useful to analyze metastable dynamics in a variety of systems, including the protein folding dynamics analyzed here.